# Forecasting Electricity Market Signals

### v4

@@ Abstract
This paper presents a generative artificial intelligence approach to probabilistic forecasting of electricity market signals, such as real-time locational marginal prices and area control error signals. Inspired by the Wiener-Kallianpur innovation representation of nonparametric time series, we propose a weak innovation autoencoder architecture and a novel deep learning algorithm that extracts the canonical independent and identically distributed innovation sequence of the time series, from which samples of future time series are generated. The validity of the proposed approach is established by proving that, under ideal training conditions, the generated samples have the same conditional probability distribution as that of the ground truth. Three applications involving highly dynamic and volatile time series in real-time market operations are considered: (i) locational marginal price forecasting for self-scheduled resources such as battery storage participants, (ii) interregional price spread forecasting for virtual bidders in interchange markets, and (iii) area control error forecasting for frequency regulations. Numerical studies based on market data from multiple independent system operators demonstrate the superior performance of the proposed generative forecaster over leading classical and modern machine learning techniques under both probabilistic and point forecasting metrics.
@@ 1 Introduction
A defining feature of generative artificial intelligence (AI) is its ability to produce artificial samples that resemble reality. In particular, generative AI learns the underlying structure of a phenomenon from examples, from which an arbitrarily large number of artificial samples having the same properties exhibited in the examples are produced. Generative AI with sophisticated neural network structure and machine learning methods have achieved remarkable performance in real-world applications unmatched by conventional techniques [1].
The classical subject of forecasting has a natural connection with generative AI through probabilistic forecasting. Given past observations, probabilistic forecasting aims to obtain the conditional probability distribution of future time series given past observations. Once such a distribution is obtained, Monte-Carlo samples of the future time series can be generated. However, forecasting conditional probability distribution faces daunting computation and sample complexities. First, nonparametric distribution forecasting is an infinite-dimensional functional estimation problem, typically requiring a finite-dimensional reduction through a histogram with a finite number of bins or quantiles with finite levels. Second, for continuously distributed random time series, there is only one realization of the future associated with the particular history on which conditional probability is defined, which makes learning the conditional distribution from data fundamentally difficult.
This paper develops a generative probabilistic forecasting (GPF) solution based on the generative AI principle that directly generates future time series samples based on current and past observations, bypassing the modeling, computational, and sample complexity challenges of forecasting the conditional probability distribution. Currently, such nonparametric GPF techniques do not exist. At the theoretical level, it is unknown what the GPF implementation architecture would be to guarantee the generation of samples with the correct conditional distribution.
@@ 1.1 Literature Review
The literature on parametric and nonparametric probabilistic forecasting of electricity market signals is vast. Our review here will focus on short-term (real-time) forecasting techniques for wholesale electricity prices and dispatch quantities such as area imbalances. To this end, we rely on earlier surveys [2, 3, 4] and pay more attention to contemporary machine learning-based techniques over the last decade or so.
@@
Because wholesale electricity prices and dispatch imbalances are endogenously determined by optimization-based market clearing, they tend to be highly volatile due to their sensitivity to binding constraints. Overall, real-time prices and dispatch quantities behave quite differently from exogenous (physical) processes such as wind, solar, and demand time series. For this reason, our review excludes the extensive literature on energy forecasting, even though many techniques also apply to price forecasting. See [5] and references therein for energy forecasting.
@@
Mainstream probabilistic forecasting techniques are from parametric or nonparametric families. Parametric forecasting predicts a parameterized conditional probability distribution of future time series variables, reducing an infinite-dimensional inference problem to one of finite-dimensional estimation. Popular approaches include autoregressive and moving average models [6, 7, 8, 9, 10, 11], Gaussian models [12], Student’s $t$-distribution [13], and others [14]. A classic benchmark with strong performance is SNARX [4], which we include in our comparison. While restricting probability distributions to be parametric results in tractable computations, it comes with the loss of accuracy from model mismatch.
@@
Nonparametric forecasting has a long history. See [2] for a review up to 1997. These techniques estimate the underlying probability distribution or its derived properties (e.g., quantile) without assuming a parametric form. Classical techniques [15] face formidable sample and computational complexities, further magnified when the underlying time series has arbitrary temporal dependencies. Quantile regression is one of the most popular techniques for forecasting electricity prices. By estimating multiple quantiles, one obtains a histogram approximation of the underlying probability density function. One well-recognized example using quantile regression for day-ahead LMP foreasting is [16]. Over the last decade, modern machine learning techniques are widely adopted to compute the conditional quantiles. Examples include Extreme Learning Machine (ELM) [17], Recurrent Neural Network (RNN) [18], and attention mechanism [19]. See [20] for a comparison study among multiple quantile regression methods. These techniques, however, are limited by the resolution of the quantile and cannot produce real-valued generative samples corresponding to the true distribution.
Most machine learning-based GPF methods involve a deep neural network (possibly with sophisticated LSTM or transformer architectures) that maps past observations and a set of exogenous random variables to samples of future time series. Distinguishing these techniques are methods used to train the neural network and ways of generating exogenous random variables. Examples include variational autoencoder (VAE) based methods [21, 22, 23, 24] and the more recent normalizing flow and denoising diffusion techniques [25, 26]. Generative adversarial network (GAN) learning has also been used [27, 28]. Missing in these methods is some form of theoretical guarantee that the generative samples should follow the desired conditional probability distribution.
Unlike point forecasting problems where the ground-truth samples are available, evaluating GPF is particularly challenging because of the lack of ground-truth distribution. Aside from standard evaluation criteria (see B), one way is to evaluate GPF by the performance of GPF-derived point forecasting techniques, since GPF can produce almost any type of point forecasts. The argument is that if a GPF method performs well, then all the GPF-derived point forecasting methods should perform well under point forecasting metrics. Thus, it is relevant to review some of the well-tested point forecasting techniques with which we compare the point forecasting techniques derived by the proposed GPF method.
One of the earliest GAN-based point forecasting techniques for real-time LMP is [29]. Very recently, the success of large language models (LLM) in natural language processing motivated its application in time series forecasting [30, 31], including the direct application of ChatGPT to price forecasting [32]. While there has been timely discussion of LLM’s capabilities and limitations in power system applications [32], there is a lack of understanding of the rationale of using highly complex LLM for LMP forecasting and the reasons for their improvement over the more conventional machine learning techniques. We include transformer-based point techniques as benchmarks for comparison in Sec. 4, given their demonstrated superior performance for natural languages and other time series. Informer [33] and Pyraformer [34] are point forecasting techniques that adopted the popular transformer model to capture the temporal dependencies. For their award-winning contributions, we include [33] and [34] in our comparative study in this paper.
@@ 1.2 Summary of Contributions
Note that the generator output $\tilde{{\bm{X}}}_{t}$ is a function of observed time series ${\bm{x}}_{0:t}$ and independently generated exogenous random vector $\tilde{\mathcal{V}}_{t}=({\bf V}_{1},\cdots,{\bf V}_{T})$ with independent and identical (IID) uniformly distributed components, making $\tilde{{\bm{X}}}_{t}$ a function of the random vector $\tilde{\mathcal{V}}_{t}$. By generating $\tilde{{\cal V}}_{t}$ from $T$ IID samples of uniform distribution, WIAE-GPF produces realizations of $\tilde{{\bm{X}}}_{t}$ following the same conditional distribution as ${\bm{X}}_{t+T}$. The formal definition of WIAE and its learning algorithm are presented in Sec. 2. The WIAE-GPF architecture and its validity are shown in Sec. 3.
@@ 4.1 Baseline Methods in Comparison
SNARX [4] Probabilistic Semiparametric AR AR Model Parameters Kernel Estimation
@@
LQRA [16] Probabilistic Nonparametric Forecasted Quantiles Ensemble method + Lasso regularization
@@
We started with SNARX [4], a classical parametric forecasting technique based on an auto-regressive moving-average (ARMA) model. An early study showed that SNARX performed the best among a list of parametric techniques [4]. For deep-learning techniques, we compared with DeepVAR [11], a multivariate generalization of the popular DeepVAR that has become a key baseline for time series forecasting in multiple applications. Temporal Latent AutoEncoder (TLAE) [21] is an autoencoder-based parametric GPF where the conditional distribution of the future time series variables is obtained through a transformed Gaussian random process with mean and variance as parameters. Once these parameters are estimated from observed realizations, Gaussian Monte Carlo samples are fed into the decoder to generate samples of the forecasted random variable. LQRA [16] builds on top of quantile regression averaging by utilizing LASSO regularization for quantile regression. It computes conditional quantiles by linear transforming a collection of point forecasts. The point forecasts are obtained by fitting an expert model. The expert model is originally proposed for day-ahead price prediction, and thus we made our adaptation by adding the real-time price for the past 5 hours to the linear autoregression.
We also included three popular forecasting techniques based on LLMs. One is the award-winning technique Informer [33]; the other is Pyraformer [34] that captures temporal dependency at multiple granularities§§§Pyraformer is the keynote presentation at The International Conference on Learning Representations (ICLR) 2022.. Pyraformer showed superior performance over a wide range of LLM-based point forecasting techniques. Specifically developed for LMP forecasting, BWGVTWe have named the algorithm by the first letters of authors’ last names [19] combines quantile regression with a transformer architecture derived for LLMs.
4.2 Evaluation Metrics
Comparing probabilistic forecasting methods is difficult due to the lack of ground truth for the underlying conditional distribution. However, because a GPF can produce arbitrarily many Monte Carlo samplesWe used $1000$ Monte-Carlo samples and sample average to obtain point estimates., it can be evaluated by all point forecasting metrics. More importantly, an ideal probabilistic forecaster that produces the correct conditional distribution will perform well under any point estimator metric under regularity conditions. Therefore, evaluating a GPF method based on a set of point-forecasting techniques is a credible way to assess its performance.

### v5

# Probabilistic Forecasting of Real-Time Electricity Market Signals

@@ Abstract
This paper introduces a generative AI approach to probabilistic forecasting of real-time electricity market signals, including locational marginal prices, interregional price spreads, and demand-supply imbalances. We present WIAE-GPF, a Weak Innovation AutoEncoder-based Generative Probabilistic Forecasting architecture that generates future samples of multivariate time series. Unlike traditional black-box models, WIAE-GPF offers interpretability through the Wiener-Kallianpur innovation representation for nonparametric time series, making it a nonparametric generalization of the Wiener/Kalman filter-based forecasting. A novel learning algorithm with structural convergence guarantees is proposed, ensuring that, under ideal training conditions, the generated forecast samples match the ground truth conditional probability distribution. Extensive tests using publicly available data from U.S. independent system operators under various point and probabilistic forecasting metrics demonstrate that WIAE-GPF consistently outperforms classical methods and cutting-edge machine learning techniques.
@@ 1 Introduction
A key feature of generative AI is its ability to produce artificial samples that closely resemble real-world data. In particular, generative AI learns the underlying structure of a phenomenon from examples, enabling it to generate an arbitrarily large number of artificial samples that exhibit the same properties as the original data. Leveraging advanced neural networks and machine learning techniques, generative AI has achieved performance in real-world applications that far surpasses conventional methods [1].
The classical field of Probabilistic Forecasting (PF) naturally aligns with generative AI. In particular, PF predicts the conditional probability distribution of the future given past time series observations. Once this distribution is known, Monte Carlo samples of the future series can be generated. However, forecasting such conditional distributions presents significant computational and sampling challenges. Specifically, nonparametric distribution forecasting is an infinite-dimensional functional estimation problem, often requiring finite-dimensional reductions, such as histograms with a finite number of bins or quantiles with a finite set of levels. Second, for continuous time series, each future realization is uniquely tied to a specific past history, making it inherently difficult to learn the true conditional distribution from data. The standard approach to PF is to assume a parametric model, reducing the infinite dimensional forecasting problem to finite dimensional prediction of distribution parameters.
This paper applies generative AI principles to derive non-parametric Generative Probabilistic Forecasting (GPF), bypassing the obstacles of modeling, computational, and sample complexities associated with forecasting conditional probability distributions. Currently, no nonparametric GPF techniques exist. By developing practical GPF solutions, we demonstrate the somewhat surprising conclusion that GPF problem is simpler and and more practical than PF.
As powerful as generative AI has become, it is often criticized as mysterious and uninterpretable, which casts doubt on whether the generated samples indeed follow the ground-truth distribution or merely appear to resemble the training data. For time series forecasting problems, it is difficult, if not impossible, to compare the similarities between training data and generated samples. No existing GPF techniques have established some level of guarantee the generative samples follow the correct conditional distributions. This paper aims to derive an interpretable GPF by making connections to classic theory of time series representation pioneered by Wiener, Kallianpur, and Rosenblatt [2, 3].
@@ 1.1 Literature Review
The literature on parametric and nonparametric probabilistic forecasting of electricity market signals is extensive. In this review, we focus on short-term (real-time) forecasting techniques for wholesale electricity prices and dispatch quantities, such as area demand-supply imbalances. Drawing on previous surveys [4, 5, 6], we place particular emphasis on machine learning-based techniques developed in the past decade.
@@
Wholesale electricity prices and dispatch imbalances are endogenously determined by optimization-based market clearing processes, making them highly volatile due to their sensitivity to binding constraints. Consequently, real-time prices and dispatch quantities exhibit behavior distinct from exogenous physical processes like wind, solar, and demand time series. For this reason, we exclude the extensive body of energy forecasting literature, although those techniques can also be applied to price forecasting as well (for more on energy forecasting, see [7] and references therein). Also excluded are forecasting techniques from the market operator’s perspective, where network parameters and offers/bids are known. See, e.g., [8] and references therein.
@@
Probabilistic forecasting methods generally fall into parametric or nonparametric categories. Parametric approaches model future time series variables by predicting a parameterized conditional probability distribution, thus reducing an infinite-dimensional inference problem to a finite-dimensional. Popular parametric methods include autoregressive and moving average models [9, 10, 11, 12, 13, 14], Gaussian models [15], Student’s t-distribution [16], and others [17]. While parametric models offer computational tractability, they often sacrifice accuracy due to model mismatches.
@@
Nonparametric forecasting has a long history (see [4] for a review up to 1997). These methods estimate the underlying probability distribution or its properties, such as quantiles, without assuming a specific parametric form. However, classical nonparametric techniques [18] face significant sample and computational challenges, especially when time series exhibit complex temporal dependencies. Quantile regression is among the most popular techniques for forecasting electricity prices. By estimating multiple quantiles, it approximates the underlying probability density function through a histogram. A well-known application of quantile regression for day-ahead LMP forecasting can be found in [19].
Over the past decade, deep learning technologies have significantly advanced point, PF, and GPF generative methods, utilizing various architectures and learning principles. Examples of these include Extreme Learning Machines (ELM) [20], Recurrent Neural Networks (RNN) [21], Variational Autoencoders (VAE) [22, 23, 24, 25], Long Short-Term Memory (LSTM) networks, diffusion models [26], Generative Adversarial Network (GAN) [27] and Large Language Models (LLMs) featuring transformers and attention mechanisms [28].
Among the deep-learning-based PF and GPF techniques, the VAE-based method [22] is the most related to the proposed WIAE-GPF approach; both share a similar forecasting architecture, but differ in the training of the autoencoders. VAE typically relies on a parametric assumption, often on a conditional Gaussian distribution for the latent process, whereas WIAE-GPF imposes the innovation constraints that requires the latent process being independent and identically distributed and uniform (IID-uniform). See Sec. 2.
LLM-based generative AI, originally developed for linguistic time series, have recently demonstrated superior performance in various applications, including electricity price forecasting [29]. The transformer architecture, with its attention mechanism, has played a pivotal role in this success. However, despite their impressive results, these non-interpretable AI techniques give an understanding of the factors that would lead to effective probabilistic forecasting. In particular, there is no theoretical guarantee that LLM-based GPF can generate samples with the correct conditional probability distribution even when the training sample size is unbounded. Additionally, comprehensive empirical studies comparing LLMs to other machine learning methods for price forecasting remain scarce. To address this, we include three award-winning LLM-based models [30, 31] in our numerical comparisons in Section X, to benchmark their performance against other techniques.
@@ 1.2 Summary of Contributions
WIAE-GPF stands out as an interpretable GPF model because of its connection to the classic Wiener/Kalman predictor under parametric Gaussian and state-space assumptions. Specifically, the Innovation Encoder in Fig. 1 functions analogously to a causal whitening filter (via spectral factorization), while the Decoder corresponds to the linear Wiener predictor. Furthermore, the encoder’s operation parallels the measurement update in Kalman filtering, and the decoder mirrors the prediction (time) update. In essence, WIAE-GPF can be viewed as a non-Gaussian nonparametric extension of the Wiener/Kalman predictor.
Note that the WIAE-GPF output $\tilde{{\bm{X}}}_{t}$ is a function of observed time series ${\bm{x}}_{0:t}$ and independently generated exogenous random vector $\tilde{\mathcal{V}}_{t}=({\bf V}_{1},\cdots,{\bf V}_{T})$ with IID uniformly distributed components, making $\tilde{{\bm{X}}}_{t}$ a function of the random vector $\tilde{\mathcal{V}}_{t}$. By generating $\tilde{{\cal V}}_{t}$ from $T$ IID samples of uniform distribution, WIAE-GPF produces realizations of $\tilde{{\bm{X}}}_{t}$ following the same conditional distribution as ${\bm{X}}_{t+T}$. The formal definition of WIAE and its learning algorithm are presented in Sec. 2. The WIAE-GPF architecture and its validity are shown in Sec. 3.
@@ 3.3 From GPF to Point and Quantile Forecasting
3.4 Evaluation Metrics
Comparing probabilistic forecasting methods is difficult due to the lack of ground truth for the underlying conditional distribution. However, because a GPF can produce arbitrarily many Monte Carlo samples§§§We used $1000$ Monte-Carlo samples and sample average to obtain point estimates., it can be evaluated by all point forecasting metrics. More importantly, an ideal probabilistic forecaster that produces the correct conditional distribution will perform well under any point estimator metric under regularity conditions. Therefore, evaluating a GPF method based on a set of point-forecasting techniques is appropriate to assess its performance. To this end, we also compared WIAE-GPF with some of the well-tested point forecasting techniques.
@@ 4.1 Baseline Methods in Comparison
For deep-learning techniques, we compared with DeepVAR [14], a multivariate generalization of the popular DeepVAR that has become a key baseline for time series forecasting in multiple applications. Temporal Latent AutoEncoder (TLAE) [22] is an autoencoder-based parametric GPF where the conditional distribution of the future time series variables is obtained by passing a conditional Gaussian process through the decoder, where the conditional mean and variance are computed by the encoder. Once these parameters are estimated from observed realizations, Gaussian Monte Carlo samples are fed into the decoder to generate samples of the forecasted random variable.
We also included three popular forecasting techniques based on LLMs. One is the award-winning technique Informer [31]; the other is Pyraformer [30] that captures temporal dependency at multiple granularitiesPyraformer is the keynote presentation at The International Conference on Learning Representations (ICLR) 2022.. Pyraformer showed superior performance over a wide range of LLM-based point forecasting techniques. Specifically developed for LMP forecasting, BWGVTWe have named the algorithm by the first letters of authors’ last names [28] combines quantile regression with a transformer architecture derived for LLMs.
@@ 4.2 LMP forecasting for Energy Market Participation
The real-time market of NYISO closes sixty minutes ahead of actual delivery, which means that the forecasting horizon needs to be longer than 60 minutes. Its real-time LMPs and load were collected every 5 minutes and day-ahead LMPs every hour. Two experiments were conducted to produce probabilistic forecasts of 60-minute ahead LMPs at the Long Island (LONGIL) using (a) the day-ahead prices and the current and past real-time LMPs at LONGIL, along with the system load up to the time of submitting the bid; (b) the neighboring NYC real-time LMP in addition to the data in (a).
Electricity prices are seasonal. We selected July 2023, October 2023, January 2024, and April 2024 to represent summer, fall, winter, and spring, respectively. Our method, along with all baseline methods in Table. 2, were tested using NYISO real-time price data at LONGIL collected during these months. For every week in the selected months, a forecastor was trained using historical data from the preceding 30-day period. The same forecastor was used to produce forecast for that week.
Fig. 4 shows the real-time LMP trajectories at both LONGIL and NYC in July 2023, along with the demand and the day-ahead LMP at LONGIL. The real-tine LMP at LONGIL exhibited the highest level of volatility for all four months. The real-time LMPs at LONGIL and NYC showed apparent spatial dependencies, while the dependency between day-ahead and real-time LMPs at LONGIL. The dependencies between load and real-time LMP were less obvious. The real-time LMPs for the other three seasons show similarities to those in summer, with fall and summer exhibiting the most volatility, while spring is the least volatile. During summer and fall, real-time LMPs frequently spiked above 1000\$/MWh due to high demand and weather-related factors. In contrast, winter and spring experienced more stable prices; though on April 29th at 7:40 PM, the real-time LMP at LONGIL surged to 6752\$/MWh before dropping back below 1000\$/MWh by 8 PM.
Table 3: Comparison of volatility between seasons
Season Summer Fall Winter Spring Spring (excl. Apr. 29th, 2024)
Standard Deviation 52.8266 68.3016 43.6466 96.8191 21.6991
The standard deviation for LMPs of each seasons are included in Table. 3. Although spring LMPs appear to have the highest volatility, the extreme outlier price (over 6500\$/MWh) on Apr. 29th, 2024 mainly contributes to the high standard deviation. Excluding the extreme price from the dataset, spring LMPs exhibited the least level of volatility. Winter volatility is higher than that of spring time (excluding Apr. 29th), though it is still lower than that of Summer and Fall.
As illustrated by Fig. 4, the real-time LMPs at LONGIL exhibit significant spikes with absolute values far exceeding the rest. These spikes are challenging to predict, and even minor prediction errors for these spikes can disproportionately impact the overall evaluation metrics. To ensure a more informative comparison among techniques, we calculated the evaluation metrics only on the “normal” LMPs, defined as those within three standard deviations from the mean.”
Real-time LMP forecasting performance: an overall summary
We analyzed the overall performance of all methods across the four seasons. WIAE-GPF consistently achieved the best overall performance on all evaluation metrics, both point and probabilistic. It ranked first in NMSE, CRPS, and CPE for nearly all cases, while it placed second in NMAE, MASE, and sMAPE in most instances. The strong performance of WIAE-GPF can be attributed to its focus on matching the conditional distribution, with a validity guarantee ensuring that the Monte Carlo samples generated have the same conditional distribution as the actual time series variable. The strong performance of WIAE-GPF compared with benchmark techniques under NMSE, NMAE, and other point forecasting metrics suggests that the WIAE-GPF approximates the true conditional probability distribution well.
The second best-performing technique overall was DeepVAR, which excelled in NMAE and MASE for summer and fall. This indicates that DeepVAR achieved accurate parameter estimation for the mean and median of the Gaussian AR model it assumes. However, its CPE and CRPS metrics are generally worse than those of WIAE-GPF, possibly due to a model mismatch between the volatile and complex conditional distribution of real-time LMP and the Gaussian AR assumption of DeepVAR.
The three LLM-based estimators (BWGVT, Pyraformer, and Informer) performed similarly, outperforming TLAE but generally falling short of WIAE-GPF and DeepVAR. Pyraformer and Informer, both point forecasters trained to minimize mean squared forecasting error, performed better under NMSE. Notably, Pyraformer achieved the best NMSE among all methods during the winter month. However, their NMAE, MASE, and sMAPE scores were generally worse than those of WIAE-GPF and DeepVAR in most cases. Informer underperformed compared to Pyraformer across the other seasons, but it did achieve the best NMAE and MASE during the spring month, when LMP exhibited substantially lower volatility. We conclude the underperformance of the LLM-based point estimators to the un-necessity of adopting attention mechanism for long-range dependency and the difficulty of training due to their large numbers of parameters. For details, see Sec. 4.5.
The LLM-based probabilistic forecasting technique BWGVT uses quantile regression to predict future distributions. Its quantile predictions were sensitive to outliers, especially compared to the GPF methods that rely on a stochastic latent process, leading to poorer point forecasting performance. BWGVT tended to predict larger intervals, as indicated by its high NCW across all cases. As a result, its CPEs were always positive, meaning that the prediction intervals it generated covered more than the nominal percentages. Consequently, both its point estimation results and CRPS were worse than those of DeepVAR and WIAE-GPF.
TLAE performed the worst across the broad. Its point evaluation metrics consistently ranked the worst over all methods, while its probabilistic evaluation metrics are slighly better. TLAE is a VAE-based auto encoder with a correlated Gaussian latent process. It generates samples of forecasts at the next time stamp by re-parameterization. For a multiple timestamp-ahead prediction, TLAE drawns sample iteratively by substituting the mean of the samples generated for the previous timestamp. This heuristic has no guarantee to match the conditional distribution of the generated samples and the conditional distribution of the future given the current observations, and it could also cause an accumulation of biases.